OpenStudy (anonymous):

Write v as a linear combination of u1,u2, and u3 if possible u1=<1,3,5> u2=<2,-1,3> u3=<-3,2,-4> v=<-1,7,2>

OpenStudy (anonymous):

i dont get it....

OpenStudy (anonymous):

find a linear combination (combination of u1 and u2 and u3) to make v

OpenStudy (anonymous):

you are going to need variables

OpenStudy (anonymous):

LOL...wat math is this?

OpenStudy (anonymous):

linear algebra

OpenStudy (anonymous):

do you know vectors?

OpenStudy (anonymous):

im in alg 2...i never heard of this..is this college? cuz im still in hs freshman year

OpenStudy (anonymous):

it's ehh algebra sort of would it help if it if i made it more simple like a(1,3,5)+b(2,-1,3)+c(-3,2,-4)=(-1,7,2)

OpenStudy (anonymous):

i dunt really under stand it, but i will still give u a medal

OpenStudy (anonymous):

lol i don't want medal teacher doesn't give answers to pretests so

OpenStudy (anonymous):

srry...dnno, ask sumbody else, like saifoo

OpenStudy (anonymous):

\[a+b-3c=-1\] \[3a-b+2c=7\] \[5a+3b-4c=2\]

OpenStudy (anonymous):

first one should be 2b

OpenStudy (anonymous):

how do u do it using matrices cuz i noe ima mess up wen using subs and elim

OpenStudy (anonymous):

hehe you can check it when you aren't supposed to use matrices that's how a lot of people on here find out the answers to long linear equations like this. if you have a graphing calculator you can put 2nd>x^-1>move to edit>ENTER> dimensions of the matrix will be rows x colums so in this case 3x4>then put the coefficients into matrix so like 1,2-3,-1]

OpenStudy (anonymous):

should look something like this